OpenStudy (anonymous):

47 lbs = ______________ kgs (give answer in decimal form)

OpenStudy (anonymous):

21.3188 kgs

OpenStudy (anonymous):

googled

OpenStudy (anonymous):

i need to show my work so I need to know how to get the answer but thanks.

OpenStudy (anonymous):

1 pound = 0.45 47 = x cross multiply 1x = 47 * 0.45 = 21.15

hero (hero):

\[1 \space\text{kg} = 2.2\space\text{lbs}\] \[47\space\text{lbs} \times \frac{\text{kg}}{2.2\space\text{lbs}} = \frac{47}{2.2}\text{kg}\]